if roots of equation (n-p)(n-q)=r, in n are l and m, then the roots of equation (n-l)(n-m)+r=0 is p and q..prove it plz

Grade:11

2 Answers

Sunil Raikwar
9 years ago
(n-p)(n-q)-r=(n-l)(n-m)
(n-p)(n-q)=(n-l)(n-m)+r
Hence p & q are the roots of the equation (n-l)(n-m)+r=0
Thanks & Regards
Sunil Raikwar
askIItians faculty
Manas Satish Bedmutha
9 years ago
If l and m are roots of 1 st equation, (n-p)(n-q) – r = (n-l)(n-m)
Hence (n-l)(n-m) + r = (n-p)(n-q) = 0 as given.
Hence, n = p or q. Hence proved